# Class-Independent Increment: An Efficient Approach for Multi-label Class-Incremental Learning

### v1

@@ III Method
Figure 2: The framework of proposed CLIN. During training, the image $x^{t}$ (In actuality, an input image can contain either one or multiple new classes, and it can also contain one, multiple, or no old classes.) undergoes feature extraction by the backbone and is then converted into patch tokens $x_{P}$. Together with a set of class-specific tokens, this forms the input for the class-level cross-attention module to generate class-level embeddings $E^{t}$. The set of class-specific tokens includes trainable tokens $\{\boldsymbol{q}_{c}\}_{c=1}^{C_{n}}$ for new classes and frozen tokens $\{\boldsymbol{q}_{c}\}_{c=1}^{C_{o}}$ for old classes, and compute $\mathcal{L}_{to}$ to maintain orthogonality. Subsequently, $E^{t}$ are separately fed into the class-independent classifier to compute $\mathcal{L}_{ce}$ and a projection layer to calculate $\mathcal{L}_{mc}$.
III-A Problem Formulation
@@
The multi-label class-incremental learning problem [11, 12] is defined as follows: suppose we have T sequences of multi-label image sets $\left\{\mathbf{D}^{1},\mathbf{D}^{2},...,\mathbf{D}^{T}\right\}$, where $\mathbf{D}^{t}=\left\{\mathbf{X}^{t},\mathbf{Z}^{t}\right\}$ is consisted of a training set $\mathbf{X}^{t}$ and a test set $\mathbf{Z}^{t}$. Each training set is defined as $\mathbf{X}^{t}=\left\{\left(\mathbf{x}_{i}^{t},\mathbf{y}_{i}^{t}\right)\right\}_{i=1}^{N^{t}}$, where $N^{t}$ represents the number of training samples, $\mathbf{x}_{i}^{t}$ is the $i$-th training sample and $\mathbf{y}^{t}_{i}\subseteq\mathbf{C}^{t}$ is a label set with $1\leq|\mathbf{y}^{t}_{i}|\leq|\mathbf{C}^{t}|$. $\mathbf{C}^{t}$ denotes the class collection at the $t$-th session and $\forall m,n~{}(m\neq n),~{}\mathbf{C}^{m}\cap\mathbf{C}^{n}=\varnothing$ [^1] The training set images from different sessions can be repeated, but the annotated labels are different in [11], i.e., $X_{m}\cap X_{n}\neq\emptyset$. In protocol [12], image repetition is strictly restricted, i.e., $X_{m}\cap X_{n}=\emptyset$.. In the MLCIL setting, a multi-label classification model is required to incrementally learn a unified classifier from a sequence of training sessions. At session $t$, only $\mathbf{X}^{t}$ is available during training and the model is evaluated on a combination of test sets $\mathbf{Z}^{1\sim t}=\mathbf{Z}^{1}\cup\cdots\cup\mathbf{Z}^{t}$ and is expected to recognize all the encountered classes $\mathbf{C}^{1\sim t}=\mathbf{C}^{1}\cup\cdots\cup\mathbf{C}^{t}$ present in the test images.
III-B Overall Framework
@@
Assuming the $t$-th incremental session, the input of image $x^{t}$ from $\mathbf{D}^{t}$ is fed into the feature extractor $f(\cdot;\theta)$ and projected into the image token $X_{P}\in\mathbb{R}^{L\times D}$, where $L$ is the token length and $D$ is the feature dimension. Following this, $X_{P}$ serves as input to the CINet and obtains class-level embeddings $E^{t}$. In the CINet, the $\gamma(\cdot;\omega)$ extends $|\mathbf{C}^{t}|$ ($C_{n}$) class-specific tokens (Orange one) to learn new classes and freeze $|\mathbf{C}^{1\sim t-1}|$ ($C_{o}$) old tokens (Blue one) to prevent forgetting of old classes. To alleviate potential conflicts between the existing tokens and trainable tokens, we orthogonally initialize these tokens and maintain orthogonality through a token orthogonal loss $\mathcal{L}_{to}$.
@@
III-C The Structure of Proposed CLIN
@@
(2) Class-independent Classifier: The input of the class-independent classifier is the class-level embeddings $E=\{e_{1},e_{2},...e_{|\mathbf{C}|}\}$, where the class-level embedding $e_{c}$ captures the features of the input image in the context of label $c$. To establish the correspondence between class labels and class-level embeddings and to ensure no interference between different categories, especially between old and new classes, we utilize $|\mathbf{C}|$ separate Feed Forward Networks (FFNs) as classifiers instead of joint classifiers. It predicts the presence of corresponding labels through binary classification.
@@
$p_{c}=\sigma(w_{c}\times e_{c}+b_{c}),\,\,c\in[1,|\mathbf{C}|]$ (4)
@@
$\mathbf{X}_{q}=[{\color[rgb]{0,0,1}\mathbf{X}_{q}^{1}},...,{\color[rgb]{0,0,1}\mathbf{X}_{q}^{t-1}},{\color[rgb]{1,.5,0}\mathbf{X}_{q}^{t}}],\,\,X_{q}^{t}=\{\boldsymbol{q}_{c}\}_{|\mathbf{C}^{1\sim t-1}|+1}^{|\mathbf{C}^{1\sim t}|}$ (5)
@@
Where the new tokens $X^{t}_{p}$ are trainable while the old tokens $X^{1\sim t-1}_{q}$ are frozen in the $t$-th training session. The frozen tokens effectively preserve the knowledge of previous classes to mitigate catastrophic forgetting while simultaneously optimizing the trainable parameters to learn new classes. Then, the patch tokens $X_{p}$ and class-specific tokens $X_{q}$ learn attention weights regarding specific embeddings through a multi-head attention mechanism, obtaining class-level embeddings $E\in\mathbb{R}^{|\mathbf{C}^{1\sim t}|\times D}$, which are then input into the classifier for classification.
@@
To learn new classes, the class-independent classifier also adds $|\mathbf{C}^{t}|$ Feed Forward Networks (FFNs). Similar to class-specific tokens, FFNs are categorized into two types: frozen FFNs $\varphi(\cdot;\phi_{1})$,…,$\varphi(\cdot;\phi_{|C^{1\sim t-1}|})$ and trainable FFNs $\varphi(\cdot;\phi_{|C^{1\sim t-1}|+1})$,…,$\varphi(\cdot;\phi_{|C^{1\sim t}|})$. These serve the purpose of maintaining old knowledge and learning new knowledge. Ultimately, the image’s classification loss $\mathcal{L}_{ce}$ is calculated based on the predicted probabilities from each FFN.
III-D The Loss Function of CLIN
@@
After introducing the structure of the CLIN method, we proceed to discuss the loss functions that optimize the CLIN method. To address the feature confusion problem, we introduce two terms in the loss function to optimize the learning of class-level embeddings and class-specific tokens: multi-label contrastive loss and token orthogonal loss. Additionally, similar to previous approaches, our loss function includes base losses, such as classification loss and distillation loss, aiming to achieve the classification task and prevent catastrophic forgetting.
@@
Multi-label Contrastive Loss: In contrast to regular supervised learning on independently and identically distributed data, continual learning faces challenges such as poor discriminative ability in non-stationary data distribution. To address this issue, regular supervised learning methods often employ techniques like metric learning [72] and contrastive learning [73]. Since image-level features are associated with multiple labels, directly applying them to multi-label tasks is challenging. For example, assuming that the image-level representation of an image containing a car must always be close to each other is unreasonable because the car is just one of many objects in these images, possibly occupying only a small part of the image. However, in our CLIN approach, we tackle this problem by using attention mechanisms to transform image features into class-level features, effectively converting the multi-label problem into multiple single-label problems. Furthermore, we propose a multi-label contrastive loss to force the model to learn diverse and discriminative features for new concepts, further addressing the feature confusion problem.
@@
Specifically, in the $t$-th session, given a small batch $N$ of images $X^{t}=\{x^{t}_{i}\in\mathbf{D}^{t}\,|\,i\in\{1,\ldots,N\}\}$ (we omit the superscript $t$ in the following text as it is unrelated to session). We first input them into the class-level cross-attention module and the projection layer, obtaining class-level embeddings in the mapped space and defining them as the set $Z=\{z_{ic}\in\mathbb{R}^{D}\,|\,i\in\{1,\ldots,N\};c\in\{1,\ldots,|\mathbf{C}|\}\}$. Similarly, we define the set of ground truth labels of the minibatch: $Y=\{y_{ic}\in\{0,1\}\,|\,i\in\{1,\ldots,N\};c\in\{1,\ldots,|\mathbf{C}|\}\}$. It’s worth noting that we consider the class-level embedding $z_{ic}$ of an image as an instance rather than the image itself; $z_{ic}$ is associated with a single true label $y_{ic}$. We further define $A=\{\{a_{k},y_{k}\}|\{z_{ic}\in Z\,and\,y_{ic}=1\}\}$ as the instance set containing class-level embeddings, where $a_{k}$ represents the $k$-th instance in $A$, $y_{k}$ represents the label of $a_{k}$, and $y_{k}=c$.
For pairs of class-level embeddings $a_{k}$ and $a_{m}$, which are randomly selected from the instance set $A$. Our inter-class objective enforces the following loss on the model:
$\mathcal{L}_{mc}=\frac{1}{|A|}\sum_{k=1}^{|A|}\sum_{m=1}^{|A|}(1-cos(a_{k},a_{m}))*B_{km}.$ (6)
@@
Token Orthogonal Loss: In the class-level cross-attention module, the class-specific tokens expand gradually based on the number of classes. To prevent interference between learning new knowledge and old knowledge while enhancing the model’s ability to learn diverse concepts and foster class-specific distinctions for retaining knowledge, we propose addressing this issue through appropriate initialization and a specific loss function.
We apply the Gram-Schmidt process to initialize new token parameters at the beginning of each new session. Certainly, at the $t$-th session, $|\mathbf{C}^{t}|$ class-specific tokens $X^{t}_{q}$ are introduced. Each new token undergoes the Gram-Schmidt process to ensure that the initialized parameters are orthogonal to the parameters of all the old tokens. This process ensures that the new class-specific tokens are linearly independent from the previously existing tokens. This helps maintain diversity and prevents interference between the representations of different classes, facilitating effective learning and discrimination.
Additionally, we employ a token orthogonal loss to maintain orthogonality between trainable tokens (new classes) and frozen tokens (old classes). In the session $t>1$, the token orthogonal loss compels the model to optimize new class tokens $X_{q}^{t}$ (Orange) to be orthogonal to previous tokens $X_{q}^{1\sim t-1}=[X_{q}^{1},X_{q}^{2},...X_{q}^{t-1}]$ (Blue). This constraint can be formulated as follows:
$\mathcal{L}_{to}=$ $\|[{\color[rgb]{0,0,1}X_{q}^{1}},...,{\color[rgb]{0,0,1}X_{q}^{t-1}}]\cdot{\color[rgb]{1,.5,0}(X_{q}^{t})}^{T}\|$ (7)
[…]
@@
$\mathcal{L}_{ce}=\frac{1}{N}\sum_{n=1}^{N}\left\{\begin{aligned} &(1-p_{c})^{\gamma+}log(p_{c}),&y_{c}=1,\\ &p_{c}^{\gamma-}log(1-p_{c}),&y_{c}=0,\end{aligned}\right.$ (8)
@@
Total Optimization Objective: For session $t~{}(t>1)$, our full optimization consists of the two portions: (1) the separation loss includes the multi-label contrastive loss $\mathcal{L}_{mc}$ and token orthogonal loss $\mathcal{L}_{to}$. (2) the base loss includes the classification loss $\mathcal{L}_{ce}$ and distillation loss $\mathcal{L}_{kd}$. We can obtain the total loss function:
$\mathcal{L}=\mathcal{L}_{ce}+\alpha*\mathcal{L}_{mc}\ +\beta*\mathcal{L}_{kd}+\lambda*\mathcal{L}_{to}.$ (10)
@@ IV Experiments
In this section, we conduct extensive experiments to validate the effectiveness of our algorithm, particularly by evaluating its performance on the MS-COCO [74] and PASCAL VOC 2007 [75] datasets using two widely recognized benchmark protocols [11, 12]. Additionally, we delve into a series of ablation studies and visualizations aimed at assessing the importance of each component, providing deeper insights into the workings of our method. The section is structured as follows: we begin by introducing the experimental setup and implementation details in Sec. IV-A. Subsequently, we present the experimental results on the two protocols in Sec. IV-B. Finally, we introduce the ablation study and provide a comprehensive analysis of our method in Sec. IV-C.
IV-A Experimental Setup and Implementation Details
@@
Dataset. MS-COCO [74] is a large-scale dataset constructed for segmentation and object detection tasks first and has been widely used for evaluating MLC. It contains 122,218 RGB images of 80 object categories, where each class has 82,081 images in the training set and 40,137 images in the validation set. PASCAL VOC [75] is a frequently used dataset for MLC and consists of $9,963$ images for 20 object categories, where 5,011 images are in the train-value set, and 4,952 images are in the test set.
Benchmark Protocols. We conduct comprehensive experiments on two popular benchmark protocols [11, 12]. The first protocol is derived from the approach outlined in [11]: arranging class names in alphabetical order and then partitioning the training set into multiple incremental sessions based on the order of categories. While different sessions may include the same images, they come with different provided labels. For the division of incremental phases, we adopt protocols commonly used in CIL [26, 24]. Specifically, for the MS-COCO benchmark, we define it as COCO-B0CX and COCO-B40CY, where B represents the number of base classes, and CX(Y) denotes the addition of X(Y) new class categories during the incremental sessions (X=10, 20, and Y=5, 10). The B0 presents a more challenging scenario. Similarly, for the VOC benchmark, we evaluate our method using the B0-C4 and B10-C2 protocols (denoted as COCO and VOC).
TABLE I: Main class-incremental results on MS-COCO dataset. Compared methods are grouped based on different buffer sizes. Buffer Size 0 means no rehearsal is required, where most SOTA CIL methods are not applicable anymore. All the input size is $224\times 224$ and all metrics are in %. All compared results are obtained from KRT [11]
@@
FT [61] Baseline 0 38.3 6.1 13.4 16.9 ($\downarrow 54.5$) 35.1 6.0 13.6 17.0 ($\downarrow 58.3$)
PODNet [23] CIL 43.7 7.2 14.1 25.6 ($\downarrow 45.8$) 44.3 6.8 13.9 24.7 ($\downarrow 50.6$)
oEWC [36] CIL 46.9 6.7 13.4 24.3 ($\downarrow 47.1$) 44.8 11.1 16.5 27.3 ($\downarrow 48.0$)
LWF [38] CIL 47.9 9.0 15.1 28.9 ($\downarrow 42.5$) 48.6 9.5 15.8 29.9 ($\downarrow 45.4$)
KRT [11] MLCIL 74.6 55.6 56.5 65.9 ($\downarrow 5.5$) 77.8 64.4 63.4 74.0 ($\downarrow 1.3$)
Ours MLCIL 78.1 61.2 61.1 71.4 ($\downarrow 0.0$) 79.0 65.2 64.9 75.3 ($\downarrow 0.0$)
TPCIL [22] CIL 5/class 63.8 20.1 21.6 50.8 ($\downarrow 21.8$) 63.1 25.3 25.1 53.1 ($\downarrow 23.0$)
PODNet [23] CIL 65.7 13.6 17.3 53.4 ($\downarrow 19.2$) 65.4 24.2 23.4 57.8 ($\downarrow 18.3$)
DER++ [24] CIL 68.1 33.3 36.7 54.6 ($\downarrow 18.0$) 69.6 41.9 43.7 59.0 ($\downarrow 17.1$)
KRT [11] MLCIL 75.8 60.0 61.0 68.3 ($\downarrow 4.3$) 78.0 66.0 65.9 74.3 ($\downarrow 1.8$)
Ours MLCIL 78.6 63.9 63.9 72.6 ($\downarrow 0.0$) 79.2 66.9 67.9 76.1 ($\downarrow 0.0$)
iCaRL [6] CIL 20/class 59.7 19.3 22.8 43.8 ($\downarrow 30.7$) 65.6 22.1 25.5 55.7 ($\downarrow 21.4$)
BiC [28] CIL 65.0 31.0 38.1 51.1 ($\downarrow 23.4$) 65.5 38.1 40.7 55.9 ($\downarrow 21.2$)
ER [25] CIL 60.3 40.6 43.6 47.2 ($\downarrow 27.3$) 68.9 58.6 61.1 61.6 ($\downarrow 15.5$)
TPCIL [22] CIL 69.4 51.7 52.8 60.6 ($\downarrow 13.9$) 72.4 60.4 62.6 66.5 ($\downarrow 10.6$)
PODNet [23] CIL 70.0 45.2 48.7 58.8 ($\downarrow 15.7$) 71.0 46.6 42.1 64.2 ($\downarrow 12.9$)
DER++ [24] CIL 72.7 45.2 48.7 63.1 ($\downarrow 11.4$) 73.6 51.5 53.5 66.3 ($\downarrow 10.8$)
KRT [11] MLCIL 76.5 63.9 64.7 70.2 ($\downarrow 4.3$) 78.3 67.9 68.9 75.2 ($\downarrow 1.9$)
Ours MLCIL 79.5 66.6 67.9 74.5 ($\downarrow 0.0$) 79.7 69.8 71.4 77.1 ($\downarrow 0.0$)
The second protocol is based on [12], where the training set is initially sorted using a random seed (1998). Subsequently, all training data is directly partitioned based on the tasks. It’s important to note that images in different sessions are entirely non-overlapping, although there may be unlabeled images. This protocol can be denoted by a unified terminology Bi-Cj, where i represents the class number to be learned in the base session, and j is the class number to be learned in each incremental session. We assess the models on the Split-COCO dataset with B40-C10 and B0-C20 protocols and on the Split-VOC dataset with B10-C5 and B0-C5 protocols (denoted as Split-COCO and Split-VOC).

### v2

@@ 3 Methodology
Figure 2: The framework of proposed CLIN. During training, the image $x^{t}$ (In actuality, an input image can contain either one or multiple new classes, and it can also contain one, multiple, or no old classes.) undergoes feature extraction by the backbone and is then converted into patch tokens $x_{P}$. Together with a set of class-specific tokens, this forms the input for the class-level cross-attention module to generate class-level embeddings $E^{t}$. The set of class-specific tokens includes trainable tokens $\{\boldsymbol{q}_{c}\}_{c=1}^{C_{n}}$ for new classes and frozen tokens $\{\boldsymbol{q}_{c}\}_{c=1}^{C_{o}}$ for old classes. Subsequently, $E^{t}$ are separately fed into the class-independent classifier to compute $\mathcal{L}_{ce}$ and a projection layer to calculate $\mathcal{L}_{mc}$.
3.1 Problem Formulation
@@
The multi-label class-incremental learning problem [5] is defined as follows: suppose we have T sequences of multi-label image sets $\left\{\mathbf{D}^{1},\mathbf{D}^{2},...,\mathbf{D}^{T}\right\}$, where $\mathbf{D}^{t}=\left\{\mathbf{X}^{t},\mathbf{Z}^{t}\right\}$ is consisted of a training set $\mathbf{X}^{t}$ and a test set $\mathbf{Z}^{t}$. Each training set is defined as $\mathbf{X}^{t}=\left\{\left(\mathbf{x}_{i}^{t},\mathbf{y}_{i}^{t}\right)\right\}_{i=1}^{N^{t}}$, where $N^{t}$ represents the number of training samples, $\mathbf{x}_{i}^{t}$ is the $i$-th training sample and $\mathbf{y}^{t}_{i}\subseteq\mathbf{C}^{t}$ is a label set with $1\leq|\mathbf{y}^{t}_{i}|\leq|\mathbf{C}^{t}|$. $\mathbf{C}^{t}$ denotes the class collection at the $t$-th session and $\forall m,n~(m\neq n),~\mathbf{C}^{m}\cap\mathbf{C}^{n}=\varnothing$ [^1] The training set images from different sessions can be repeated, but the annotated labels are different in [5], i.e., $X_{m}\cap X_{n}\neq\emptyset$. In protocol [33], image repetition is strictly restricted, i.e., $X_{m}\cap X_{n}=\emptyset$.. In the MLCIL setting, a multi-label classification model is required to incrementally learn a unified classifier from a sequence of training sessions. At session $t$, only $\mathbf{X}^{t}$ is available during training and the model is evaluated on a combination of test sets $\mathbf{Z}^{1\sim t}=\mathbf{Z}^{1}\cup\cdots\cup\mathbf{Z}^{t}$ and is expected to recognize all the encountered classes $\mathbf{C}^{1\sim t}=\mathbf{C}^{1}\cup\cdots\cup\mathbf{C}^{t}$ present in the test images.
3.2 Overall Framework
@@
Assuming the $t$-th incremental session, the input of image $x^{t}$ from $\mathbf{D}^{t}$ is fed into the feature extractor $f(\cdot;\theta)$ and projected into the image token $X_{P}\in\mathbb{R}^{L\times D}$, where $L$ is the token length and $D$ is the feature dimension. Following this, $X_{P}$ serves as input to the CINet and obtains class-level embeddings $E^{t}$. In the CINet, the $\gamma(\cdot;\omega)$ extends $|\mathbf{C}^{t}|$ ($C_{n}$) class-specific tokens (Orange one) to learn new classes and freeze $|\mathbf{C}^{1\sim t-1}|$ ($C_{o}$) old tokens (Blue one) to prevent forgetting of old classes.
@@
3.3 The Structure of Proposed CLIN
@@
(2) Class-independent Classifier: The input of the class-independent classifier is the class-level embeddings $E=\{e_{1},e_{2},...e_{|\mathbf{C}|}\}$, where the class-level embedding $e_{c}$ captures the features of the input image in the context of label $c$. To establish the correspondence between class labels and class-level embeddings, and to ensure no interference between different categories, especially between old and new classes, we utilize $|\mathbf{C}|$ separate Feed Forward Networks (FFNs) as classifiers instead of joint classifier. It predicts the presence of corresponding labels through binary classification.
@@
$p_{c}=\sigma(w_{c}\times e_{c}+b_{c}),\,\,c\in[1,|\mathbf{C}|],$ (4)
@@
$\mathbf{X}_{q}=[{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{X}_{q}^{1}},...,{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{X}_{q}^{t-1}},{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\mathbf{X}_{q}^{t}}],\,\,X_{q}^{t}=\{\boldsymbol{q}_{c}\}_{|\mathbf{C}^{1\sim t-1}|+1}^{|\mathbf{C}^{1\sim t}|},$ (5)
@@
where the new tokens $X^{t}_{p}$ are trainable while the old tokens $X^{1\sim t-1}_{q}$ are frozen in the $t$-th training session. The frozen tokens effectively preserve the knowledge of previous classes to mitigate catastrophic forgetting, while simultaneously optimizing the trainable parameters to learn new classes. Then, the patch tokens $X_{p}$ and class-specific tokens $X_{q}$ learn attention weights regarding specific embeddings through a multi-head attention mechanism, obtaining class-level embeddings $E\in\mathbb{R}^{|\mathbf{C}^{1\sim t}|\times D}$, which are then input into the classifier for classification.
@@
To learn new classes, the class-independent classifier also adds $|\mathbf{C}^{t}|$ Feed Forward Networks (FFNs). Similar to class-specific tokens, FFNs are categorized into two types: frozen FFNs $\varphi(\cdot;\phi_{1})$$\varphi(\cdot;\phi_{|C^{1\sim t-1}|})$ and trainable FFNs $\varphi(\cdot;\phi_{|C^{1\sim t-1}|+1})$$\varphi(\cdot;\phi_{|C^{1\sim t}|})$. These serve the purpose of maintaining the old knowledge and learning new knowledge. Ultimately, the image’s classification loss $\mathcal{L}_{ce}$ is calculated based on the predicted probabilities from each FFN.
3.4 The Loss Function of CLIN
@@
After introducing the structure of the CLIN method, we proceed to discuss the loss functions that optimize the CLIN method. To address the feature confusion problem, we introduce a novel term in the loss function to optimize the learning of class-level embeddings: multi-label contrastive loss. Additionally, similar to previous approaches, our loss function includes base losses such as classification loss and distillation loss, aiming to achieve the classification task and prevent catastrophic forgetting.
@@
Multi-label Contrastive Loss: In contrast to regular supervised learning on independently and identically distributed data, continual learning faces challenges such as poor discriminative ability in non-stationary data distribution. To address this issue, regular supervised learning methods often employ techniques like metric learning [36] and contrastive learning [3]. Since image-level features are associated with multiple labels, directly applying them to multi-label tasks is challenging. For example, assuming that the image-level repre sentation of an image containing a car must always be close to each other is unreasonable because the car is just one of many objects in these images, possibly occupying only a small part of the image. However, in our CLIN approach, we tackle this problem by using attention mechanisms to transform image features into class-level features, effectively converting the multi-label problem into multiple single-label problems. Furthermore, we propose a multi-label contrastive loss to enforce the model to learn diverse and discriminative features for new concepts, further addressing the feature confusion problem.
Figure 3: Detailed illustration of the Multi-label Contrastive Loss ($\mathcal{L}_{mc}$). $\mathcal{L}_{mc}$ is enforced on positive instances to enhance feature discrimination across new and old concepts.
@@
Specifically, in the $t$-th session, given a small batch $N$ of images $X^{t}=\{x^{t}_{i}\in\mathbf{D}^{t}\,|\,i\in\{1,\ldots,N\}\}$ (we omit the superscript $t$ in the following text as it is unrelated to session). We first input them into the class-level cross-attention module and the projection layer, obtaining class-level embeddings in the mapped space and defining them as the set $Z=\{z_{ic}\in\mathbb{R}^{D}\,|\,i\in\{1,\ldots,N\};c\in\{1,\ldots,|\mathbf{C}|\}\}$. Similarly, we define the set of ground truth labels of the minibatch: $Y=\{y_{ic}\in\{0,1\}\,|\,i\in\{1,\ldots,N\};c\in\{1,\ldots,|\mathbf{C}|\}\}$. It’s worth noting that we consider the class-level embedding $z_{ic}$ of an image as an instance rather than the image itself, $z_{ic}$ is associated with a single true label $y_{ic}$. We further define $A=\{\{a_{k},y_{k}\}|\{z_{ic}\in Z\,and\,y_{ic}=1\}\}$ as the instance set containing class-level embeddings, where $a_{k}$ represents the $k$-th instance in $A$, $y_{k}$ represents the label of $a_{k}$, and $y_{k}=c$.
For pairs of class-level embeddings $a_{k}$ and $a_{m}$ which are randomly selected from the instance set $A$. Our inter-class objective enforces the following loss on the model:
$\mathcal{L}_{mc}=\frac{1}{|A|}\sum_{k=1}^{|A|}\sum_{m=1}^{|A|}(1-cos(a_{k},a_{m}))*B_{km},$ (6)
@@
$\mathcal{L}_{ce}=\frac{1}{N}\sum_{n=1}^{N}\left\{\begin{aligned} &(1-p_{c})^{\gamma+}log(p_{c}),&y_{c}=1,\\ &p_{c}^{\gamma-}log(1-p_{c}),&y_{c}=0.\end{aligned}\right$. (7)
@@
Total Optimization Objective: For session t(t>1)t~(t>1), our full optimization consists of the two portions: (1) the separation loss includes the multi-label contrastive loss $\mathcal{L}_{mc}$. (2) the base loss includes the classification loss $\mathcal{L}_{ce}$ and distillation loss $\mathcal{L}_{kd}$. We can obtain the total loss function:
$\mathcal{L}=\mathcal{L}_{ce}+\alpha*\mathcal{L}_{mc}\ +\beta*\mathcal{L}_{kd},$ (9)
@@ 4 Experiment
In this section, we conduct extensive experiments to validate the effectiveness of our algorithm, particularly by evaluating its performance on the MS-COCO [20] and PASCAL VOC 2007 [12] datasets using two widely recognized benchmark protocols [5, 33]. Additionally, we delve into a series of ablation studies and visualization aimed at assessing the importance of each component, providing deeper insights into the workings of our method. The section is structured as follows: we begin by introducing the experimental setup and implementation details in Sec. 4.1. Subsequently, we present the experimental results on the two protocols in Sec. 4.2. Finally, we introduce the ablation study and provide a comprehensive analysis of our method in Sec. 4.3.
4.1 Experimental Setup and Implementation Details
@@
Dataset. MS-COCO [20] is a large-scale dataset constructed for segmentation and object detection tasks first and has been widely used for evaluating MLC. It contains 122,218 RGB images of 80 object categories, where each class has 82,081 images in the training set and 40,137 images in the validation set. PASCAL VOC [12] is a frequently used dataset for MLC and consists of $9,963$ images for 20 object categories, where 5,011 images are in the train-value set, and 4,952 images are in the test set. Our experimental results are reported as the averages over five runs with different random seeds, with variations within ±0.3%.
Benchmark Protocols. We conduct comprehensive experiments under two standard benchmark protocols and an additional challenging setting. 1). Standard COCO Protocols. Following the protocol in [5], class names are arranged alphabetically and the training set is partitioned into multiple incremental sessions based on category order. Although different sessions may include the same images, the provided labels differ across increments. For incremental phase division, we adopt settings commonly used in CIL [14, 2]. Specifically, on the MS-COCO dataset, we evaluate our method under the B0-C10 and B40-C10 protocols, where $B$ denotes the number of base classes and $C10$ indicates the addition of 10 new classes per incremental session. 2). VOC Protocols. On the PASCAL VOC benchmark, we follow a similar incremental setup, evaluating our method under the B0-C4 and B10-C2 protocols. Here, $B$ and $C$ denote the number of base and incremental classes, respectively. 3). Challenging COCO Protocols. To further assess robustness under more demanding conditions, we adopt stricter configurations inspired by recent work [11], which introduce either more incremental stages or fewer initial classes. Specifically, on MS-COCO, we evaluate under the B20-C4 and B0-C5 settings, providing a more rigorous benchmark for incremental learning performance.
Method Buffer Size MS-COCO B20-C4 MS-COCO B0-C5
Last Acc Avg.Acc Last Acc Avg.Acc
mAP CF1 OF1 mAP mAP CF1 OF1 mAP
Joint - 81.8 76.4 79.4 - 81.8 76.4 79.4
FT [26] 19.4 10.9 13.4 36.5 22.5 15.0 23.6 48.1
ER [27] 5/class 41.9 32.9 29.8 53.0 40.1 32.9 32.3 54.6
PODNet [6] 58.4 44.0 39.1 67.7 58.2 45.1 40.8 67.2
oEWC [30] 58.9 45.7 39.3 68.4 59.09 46.4 44.9 69.0
DER++ [2] 57.3 41.4 35.5 65.5 57.9 43.6 39.2 68.2
LwF [18] 0 34.6 17.3 31.8 55.4 50.6 36.3 41.1 66.2
KRT [5] 45.2 17.6 33.0 64.0 44.5 22.6 37.5 63.1
CSC [10] 60.6 44.5 43.0 69.8 63.4 50.7 50.1 71.1
RebLL [11] 60.1 51.3 49.2 69.2 63.5 53.5 51.9 71.7
Ours 68.4 67.2 62.9 73.5 67.9 66.2 62.7 73.3
Table 1: Experimental results on MS-COCO dataset under the challenge protocols (B20-C4 and B0-C5).
@@
FT [26] Baseline 0 38.3 6.1 13.4 16.9 ($\downarrow 55.0$) 35.1 6.0 13.6 17.0 ($\downarrow 58.8$)
PODNet [6] CIL 43.7 7.2 14.1 25.6 ($\downarrow 46.3$) 44.3 6.8 13.9 24.7 ($\downarrow 51.1$)
oEWC [30] CIL 46.9 6.7 13.4 24.3 ($\downarrow 47.6$) 44.8 11.1 16.5 27.3 ($\downarrow 48.5$)
LWF [18] CIL 47.9 9.0 15.1 28.9 ($\downarrow 43.0$) 48.6 9.5 15.8 29.9 ($\downarrow 45.9$)
AGCN [8] MLCIL 72.4 53.9 56.6 61.4($\downarrow 10.5$) 73.9 58.7 59.9 69.1($\downarrow 16.7$)
KRT [5] MLCIL 74.6 55.6 56.5 65.9 ($\downarrow 6.0$) 77.8 64.4 63.4 74.0 ($\downarrow 1.8$)
CSC [10] MLCIL 78.0 64.9 66.8 72.8($\uparrow 0.9$) 78.2 65.7 67.0 75.0($\downarrow 0.8$)
Ours MLCIL 78.6 61.6 61.5 71.9 ($\downarrow 0.0$) 79.3 65.6 65.4 75.8 ($\downarrow 0.0$)
iCaRL [25] CIL 20/class 59.7 19.3 22.8 43.8 ($\downarrow 31.2$) 65.6 22.1 25.5 55.7 ($\downarrow 22.0$)
BiC [42] CIL 65.0 31.0 38.1 51.1 ($\downarrow 23.9$) 65.5 38.1 40.7 55.9 ($\downarrow 21.8$)
ER [27] CIL 60.3 40.6 43.6 47.2 ($\downarrow 27.8$) 68.9 58.6 61.1 61.6 ($\downarrow 16.1$)
TPCIL [35] CIL 69.4 51.7 52.8 60.6 ($\downarrow 14.4$) 72.4 60.4 62.6 66.5 ($\downarrow 11.5$)
PODNet [6] CIL 70.0 45.2 48.7 58.8 ($\downarrow 16.2$) 71.0 46.6 42.1 64.2 ($\downarrow 13.5$)
DER++ [2] CIL 72.7 45.2 48.7 63.1 ($\downarrow 11.9$) 73.6 51.5 53.5 66.3 ($\downarrow 11.4$)
KRT [5] MLCIL 76.5 63.9 64.7 70.2 ($\downarrow 4.8$) 78.3 67.9 68.9 75.2 ($\downarrow 2.5$)
CSC [10] MLCIL 79.6 67.8 68.6 74.8($\downarrow 0.2$) 78.7 68.2 69.7 76.0($\downarrow 1.7$)
Ours MLCIL 79.8 67.0 68.1 75.0 ($\downarrow 0.0$) 80.0 70.2 71.9 77.7 ($\downarrow 0.0$)
Table 2: Main class-incremental results on MS-COCO dataset. Compared methods are grouped based on different buffer sizes. Buffer Size 0 means no rehearsal is required, where most SOTA CIL methods are not applicable anymore. All the input size is $224\times 224$ and all metrics are in %. All compared results are obtained from KRT [5].